OpenStudy (anonymous):

what is the reference angle for 5π/3 ?

OpenStudy (ilovecake):

π/3 3 =6π - π 3 = 6π 3 - π 3= 2π - π 3 the reference number for 5π/3 is π/3 and the terminal point of 5π/3 is in Quadrant IV. EXAMPLE: Find the reference angle for (a) θ = 17π 6 (b) θ = 13π 7 Solution: (a) We have 17π 6 =18π - π 6 = 18π 6 - π 6= 3π - π 6 , therefore θ = π 6 (Quadrant II).

OpenStudy (anonymous):

\[\frac{5* 180 {}^{\circ}}{3}=300 {}^{\circ}=-60 {}^{\circ} \]